OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 4 and a6 = 34

Parth (parthkohli):

Firstly, the common difference is 5. You can see that we are adding 5 again and again. Secondly, to find out the nth term, we must use this formula: \(\Large \color{MidnightBlue}{\Rightarrow a_{n} = a_{1} + (n - 1)d }\)

Parth (parthkohli):

d is the common difference. a1 is the first term. an is the nth term of the arithmetic sequence.

OpenStudy (apoorvk):

so the first term = a = 4 ...(i) 6th term = a + 5d = 34 ...(ii) from 'i' and 'ii', d=6 now nth term of an AP is --> a + (n-1)d. a=4, d= 6, and n=32, since you need the 32nd term. substitute and evaluate.

OpenStudy (anonymous):

a6 = a1+5d => 34 = 4+5d => d=6 and a32 = a1+31d = 4+31*6 = 190

Parth (parthkohli):

Oh oh... I got the common difference wrong.

Parth (parthkohli):

Did I or did you??

OpenStudy (anonymous):

so its 190?

OpenStudy (anonymous):

yup

Parth (parthkohli):

Yes, the common difference is 6.

OpenStudy (anonymous):

sweet! okayy well how do i do this one? What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = –67

OpenStudy (apoorvk):

same way. it's a new AP. so, just the values of the first term, and common difference would change. form equations, and find out the 'a', 'd'. your n=35

OpenStudy (anonymous):

a17 =a1+16d => -67 = 13 + 16d => d =-5 then a35 = a1 + 34d = 13 + 34*(-5) =-157

OpenStudy (anonymous):

ahhh your the best thank you so much!